# Magic Mirror on the Wall, How to Benchmark Quantum Error Correction Codes, Overall ?

### v3

@@ Abstract
Quantum Error Correction Codes (QECCs) are pivotal in advancing quantum computing by protecting quantum states against the adverse effects of noise and errors. With a variety of QECCs developed, including new developments and modifications of existing ones, selecting an appropriate QECC tailored to specific conditions is crucial. Despite significant improvements in the field of QECCs, a unified methodology for evaluating them on a consistent basis has remained elusive. Addressing this gap, this paper presents the first benchmarking framework for QECCs, introducing a set of universal parameters. By evaluating eight prominent QECCs, we propose a comprehensive suite of eight parameters for their analysis. Our methodology establishes a universal benchmarking approach and highlights the complexity of quantum error correction, indicating that the choice of a QECC depends on the unique requirements and limitations of each scenario. Furthermore, we develop a systematic strategy for selecting QECCs that adapts to the specific requirements of a given scenario, facilitating a tailored approach to quantum error correction. Additionally, we introduce a novel QECC recommendation tool that assesses the characteristics of a given scenario provided by the user, subsequently recommending a spectrum of QECCs from most to least suitable, along with the maximum achievable distance for each code. This tool is designed to be adaptable, allowing for the inclusion of new QECCs and the modification of their parameters with minimal effort, ensuring its relevance in the evolving landscape of quantum computing.
@@
Quantum error correction codes (QECCs), Benchmarking, Parameters.I Introduction
Quantum computing harnesses the principles of quantum mechanics to execute tasks that are beyond the capabilities of classical computing. This advanced technology finds applications in various fields, including molecule simulation for drug development, financial optimization enhancements, machine learning acceleration, improvement of optimization tasks, and revolutionary changes in supply chain management, among others [1, 2, 3, 4, 5]. However, commercialization of these quantum computing developments incurs challenges such as qubit stability and quantum noise [6, 7]. Quantum Error Correction Codes (QECCs) are crucial for achieving fault-tolerant quantum computing in the face of unavoidable qubit noise [8, 9]. Traditional error correction methods [10] encounter unique challenges in the quantum realm, primarily due to constraints like the no-cloning theorem [11] and the wave-function collapse that occurs during qubit measurement [12]. Research in this field has led to the creation of various quantum codes, such as the five-qubit, Bacon-Shor, topological, surface, color, and heavy-hexagon code [13, 14, 15, 16, 17, 18].
@@ I-A Motivation
The landscape of quantum computing is rich with an array of QECCs, each heralding its unique set of advantages and compromises. From the foundational Shor and Steane codes to the cutting-edge realms of topological codes, the diversity of available QECCs presents a pivotal question: ”Which code is most suitable for a specific quantum computing scenario?” This query is at the heart of advancing quantum computing technologies and underscores the pressing need for a comprehensive and standardized benchmarking methodology. Despite the wealth of research assessing various QECCs, the absence of a unified framework for objective and comparative analysis across different codes stands as a significant gap. Such comparative insight is indispensable not only for the ongoing research within the quantum computing sphere but also for steering the direction of future QECC development.
Navigating through the myriad studies to deduce the optimal QECC for a particular scenario poses a formidable challenge, compounded further when multiple scenarios are considered. The complexity of determining the most appropriate QECC—factoring in the types of qubits, error rates, and other critical parameters—demands a systematic selection strategy. The necessity for a framework that can swiftly identify a suitable QECC based on specific scenario requirements is not just a matter of convenience but a crucial step forward in the practical implementation and advancement of quantum computing technologies.
@@ I-B Contribution
To our knowledge, this research marks the inaugural effort to establish a unified benchmarking methodology for QECCs. Our objective is to discern how different QECCs are ideally suited for various scenarios and to devise a systematic strategy for the appropriate selection of QECCs. We aim to provide a comprehensive framework to facilitate this process. The contributions of this study are manifold and can be outlined as follows: Comprehensive Analysis of QECCs: We conduct an in-depth analysis of eight leading QECCs utilizing eight distinct parameters: qubit overhead, threshold, type of error protection, decoding complexity, types of transversal gates, scalability, the range of qubit types it has been realized on, and overall complexity. This enables a detailed comparative overview across different codes. Scenario-dependent QECC Suitability: Our findings emphasize that the selection of the most suitable QECC is contingent upon the specific scenario of its intended application, rather than a theoretical hierarchy of QECC efficacy. We demonstrate this through the use of radar charts, showcasing the trade-offs involved in choosing different QECCs, including variations within a single QECC type. Universal QECC Selection Strategy: We introduce a universal, systematic approach for selecting a QECC tailored to a given scenario. This strategy is designed to guide users in identifying the QECC that best matches their specific requirements, considering the multifaceted nature of quantum computing scenarios. An Adaptive QECC Recommendation Tool: A pivotal contribution of our work is the development of a first of its kind, dynamic tool that assesses user-provided scenario parameters and recommends a list of QECCs, ranked from most to least suitable, along with the maximum distance achievable for each code. This tool is designed to be evergreen, allowing for the inclusion of additional QECCs and the swift modification of their parameters without necessitating significant overhauls to the system. This ensures the framework’s adaptability and relevance in the rapidly evolving quantum computing landscape. For instance, should a QECC that has not been previously realized on trapped-ion qubits later achieve this milestone, the framework can easily accommodate this new information.
@@ I-C Paper Structure
Section II provides a brief introduction to QECCs and an extensive examination of the eight QECCs featured in this research. Section III details the eight benchmarking parameters, selection criteria for codes and parameters, and the comparative analysis method. Section IV offers a thorough analysis of the eight QECCs against the selected parameters. Section V discusses a simplified code selection strategy and introduces our QECC recommendation tool. The paper concludes in Section VI.
@@ II-B Descriptions of QECCs Under Study
Shor Code: The Shor Code [24], developed by Peter Shor in 1995, marks a significant advancement in quantum error correction by being the first to correct arbitrary quantum errors, including both bit-flip and phase-flip errors, in a qubit. It encodes one qubit of information into nine physical qubits through a two-layer structure: the first layer addresses bit-flip errors by tripling the qubit, and the second layer corrects phase-flip errors by further encoding each of these into three qubits. This innovative approach allows the Shor Code to correct any single-qubit error across its nine qubits. It is described as a degenerate code, where different error types can affect the codewords similarly. The two basis states of the code are represented as and , intricately constructed from the superposition and entanglement of the nine qubits. The two basis states for the code are defined as: and . This structure ensures that the Shor Code corrects one X error per trio of qubits and a single Z error across nine qubits, making it a comprehensive quantum error-correcting code. X errors are managed by a correction circuit for each trio, identifying and fixing bit-flips. Z errors are addressed by comparing sign differences between trios with CNOT gates. Its degeneracy allows it to correct a phase flip in any qubit of a trio effectively. Although capable of correcting three bit flips in separate trios, it’s primarily a single-error corrector due to its limitations with multiple errors. Fig. 2 shows the Shor Code’s encoding, error introduction, and correction process.
@@
Steane Code: The Steane Code [33], also known as the 7-qubit code, was introduced by Andrew Steane in 1996 as a part of the Calderbank-Shor-Steane (CSS) [25, 26] family of quantum error-correcting codes. It represents a significant advancement in the field of quantum error correction, offering a symmetrical approach to correcting both bit-flip and phase-flip errors. The Steane Code maps one logical qubit to seven physical ones, drawing from the [7,4,3] Hamming code to correct single-bit errors and extending it to fix both bit-flip and phase-flip errors in qubits. Its CSS construction enables independent correction of these errors, streamlining the correction process and simplifying the required quantum circuits. The logical zero and one states in the Steane Code are defined using superpositions of the Hamming code’s even and odd weight codewords, respectively. The Steane Code employs syndrome measurements for error detection, which uniquely indicate the presence and type of error without collapsing the quantum state. This code is comprised of 6 stabilizer generators: , , , , and . Due to this structure, the Steane Code can independently correct X and Z errors using separate circuits. This dual capability is central to its functionality, allowing for the efficient correction of single-qubit errors in quantum systems. The logical and states of this code are: and ; where . Fig. 3 depicts the error-detecting circuit of the 7-qubit Steane code.
@@
Toric Code: The Toric Code [16], introduced by Alexei Kitaev, represents a novel approach to quantum error correction by leveraging topological properties. Its use of a two-dimensional lattice structure is unique, making it resistant to a broad class of local errors. In the Toric Code, qubits are placed on the edges of a two-dimensional lattice on a torus. The topology of the lattice defines logical qubits and error correction is achieved by measuring vertex and plaquette operators. These operators are defined as: . In these expressions, and represent Pauli X and Z operators applied to the qubits. The products are taken over qubits around a vertex for vertex operators, and around a plaquette for plaquette operators. Errors in the Toric Code manifest as excitations in this lattice structure. Bit-flip errors disrupt vertex operators, while phase-flip errors disrupt plaquette operators. The code detects errors by checking for these excitations. The topological nature of the code allows for the identification and correction of errors based on the collective state of multiple qubits, rather than individual qubit states. Fig. 4 represents the toric code and the acting of the X and Z stabilizers on its unique lattice structure.
@@
Surface Code: Surface Codes [34] are a class of quantum error-correcting codes that extend the principles of the Toric Code to a planar geometry, meaning it is not modeled around a torus. They are known for their high threshold error rate, making them a promising candidate for scalable quantum computation. Surface codes, similar to toric codes in stabilizers and error detection, have two main layouts: unrotated [28] and rotated [35]. Unrotated surface codes use a square lattice, with data qubits on square edges, Z-stabilizers on surfaces, and X-stabilizers at vertices. Rotated surface codes feature a tilted lattice, with X and Z stabilizers alternating in a chessboard pattern. The difference in lattice arrangement impacts their error-correction efficiency, with rotated codes generally preferred for long-distance QEC due to a higher error threshold and simpler implementation. Error detection in Surface Codes is accomplished through the measurement of these stabilizer operators. A non-trivial measurement outcome (eigenvalue -1) indicates an error. The planar nature of Surface Codes facilitates the implementation of fault-tolerant gate operations and scalable quantum computing architectures. Fig. 5 shows the two different layouts of a surface code: unrotated and rotated.
@@
Heavy Hexagon Code: Heavy hexagon codes [39] leverage a hexagonal lattice enhanced with strategically placed ’heavy’ qubits, improving error correction efficiency through increased qubit connectivity. This design boosts error tolerance, crucial for precise syndrome measurements in identifying and correcting bit-flip (X) and phase-flip (Z) errors. The mechanism ensures the integrity of quantum information by accurately pinpointing and fixing errors. This lattice stabilizer subsystem code synthesizes elements from both Bacon-Shor and surface-code frameworks, and encodes a single logical qubit into physical qubits where denotes the code distance. This lattice is characterized by its sparse connectivity, typically not exceeding three connections per qubit, making it a robust candidate for systems with fixed-frequency transmon qubits that are vulnerable to frequency collision errors. On this lattice, data qubits and ancillas are strategically distributed along the hexagonal vertices and edges. A designated subset of these ancilla qubits serves as flag qubits, crucial for the detection of high-weight errors that may result from a limited number of fault events. The stabilizers tasked with the detection of X-type errors are constituted by the product of weight-two Z-type gauge operators, which collectively form the surface code stabilizers. Meanwhile, the X-type stabilizers, reminiscent of columnar operators in the Bacon-Shor code, are deduced by aggregating products of weight-four and weight-two X-type gauge operators. Fig. 8 represents the layout and stabilizer circuits of a heavy-hexagon code.
@@ III-A Explanation of Benchmarking Parameters
Qubit Overhead: Qubit overhead is a fundamental parameter in assessing the efficiency of a QECC. It refers to the number of physical qubits needed to encode a logical qubit; lower overhead is preferable as it implies fewer resources for error correction. However, this must be balanced against the code’s ability to correct errors, which often requires more qubits for increased distance. Therefore, analyzing the trade-off between qubit overhead and error correction capabilities is crucial in evaluating the practicality of a QECC. Error Threshold: The error threshold is a critical benchmark for the robustness of a QECC as it represents the maximum physical error rate under which a QECC can effectively correct errors. A higher threshold indicates a more fault-tolerant code, making it more desirable for practical quantum computing where error rates are non-negligible. Error Protection: Error protection capability is the measure of a QECC’s ability to identify and correct a range of quantum errors, including bit flips, phase flips, and their combinations (Pauli errors). The ultimate goal is to design QECCs capable of detecting and correcting all possible quantum errors. A QECC’s versatility in handling diverse error types is a significant indicator of its effectiveness and practical utility in quantum computing environments. Decoding: The availability and efficiency of decoding algorithms are essential for the practical implementation of QECCs. Decoding is the process of interpreting syndrome measurements to identify and correct errors. The more decoding algorithms a QECC is compatible with, the more flexible and adaptable it is. The performance of a QECC is heavily reliant on the efficiency of these algorithms, which should be fast and accurate to maintain the integrity of quantum information. Transversal Gates: Transversal gates play a crucial role in the implementation of quantum error correction. They facilitate operations on encoded qubits without decoding them, thereby preserving the error correction capability. The complexity and type of transversal gates that a QECC supports can significantly affect its functionality and integration with quantum circuits. Advanced techniques like lattice surgery, while complex, offer enhanced capabilities in error correction and logical operations, marking an important consideration in QECC design. Scalability: Scalability is an essential factor for the practical application of QECCs in large-scale quantum computing. It refers to a QECC’s ability to maintain its error correction effectiveness as the system size increases. Scalability is crucial for the advancement of quantum computing, as it determines whether a QECC can be effectively used in more complex and realistic quantum systems. Realization: The Realization of a QECC across different qubit types is a testament to its adaptability and practical relevance. The versatility of a QECC in being implemented with various qubit technologies (such as superconducting qubits, trapped-ions, etc.) is vital for its applicability in diverse quantum computing architectures. This parameter is a strong indicator of a QECC’s potential for widespread adoption in the quantum computing field. Complexity: The Structural and Operational Complexity of a QECC impacts its implementability and efficiency. This encompasses the intricacy of its encoding, error detection, and correction processes. A less complex QECC might be easier to implement but may offer limited error correction capabilities. Conversely, a more complex QECC might provide robust error correction but at the cost of increased resource demands and operational challenges. Balancing complexity with effectiveness and resource efficiency is key in the design and selection of QECCs for practical quantum computing.

### v4

@@ Abstract
Quantum Error Correction Codes (QECCs) are pivotal in advancing quantum computing by protecting quantum states against the adverse effects of noise and errors. With a variety of QECCs developed, including new developments and modifications of existing ones, selecting an appropriate QECC tailored to specific conditions is crucial. Despite significant improvements in the field of QECCs, a unified methodology for evaluating them consistently has remained elusive. This paper addresses this gap by introducing a set of eight universal parameters and evaluating nine prominent QECCs for these parameters. We establish a universal benchmarking approach and highlight the complexity of quantum error correction, indicating that the choice of a QECC depends on each scenario’s unique requirements and limitations. Furthermore, we develop a systematic strategy for selecting QECCs that adapts to the specific requirements of a given scenario, facilitating a tailored approach to quantum error correction. Additionally, we introduce a QECC recommendation tool that assesses the characteristics of a given scenario provided by the user, subsequently recommending a spectrum of QECCs from most to least suitable, along with the maximum achievable distance for each code. This tool is designed to be adaptable, allowing for the inclusion of new QECCs and the modification of their parameters with minimal effort, ensuring its relevance in the evolving landscape of quantum computing.
@@
Quantum error correction codes (QECCs), Benchmarking, Parameters, QECC variations.I Introduction
Quantum computing harnesses the principles of quantum mechanics to execute tasks that are beyond the capabilities of classical computing. This advanced technology finds applications in various fields, including molecule simulation for drug development, financial optimization enhancements, machine learning acceleration, improvement of optimization tasks, and revolutionary changes in supply chain management, among others [1, 2, 3, 4, 5, 6]. However, commercialization of these quantum computing developments incurs challenges such as qubit stability and quantum noise [7, 8]. Quantum Error Correction Codes (QECCs) are crucial for achieving fault-tolerant quantum computing in the face of unavoidable qubit noise [9, 10]. Traditional error correction methods [11] encounter unique challenges in the quantum realm, primarily due to constraints like the no-cloning theorem [12] and the wave-function collapse that occurs during qubit measurement [13]. Research in this field has led to the creation of various quantum codes, such as the five-qubit, Bacon-Shor, topological, surface, color, and heavy-hexagon code [14, 15, 16, 17, 18, 19, 20].
@@ I-A Motivation
The landscape of quantum computing is rich with an array of QECCs, each heralding its unique set of advantages and compromises. From the foundational Shor and Steane codes to the cutting-edge realms of topological codes, the diversity of available QECCs presents a pivotal question: “Which code is most suitable for a specific quantum computing scenario?” This query is at the heart of advancing quantum computing technologies and underscores the pressing need for a comprehensive and standardized benchmarking methodology. Despite the wealth of research assessing various QECCs, the absence of a unified framework for objective and comparative analysis across different codes stands as a significant gap. Such comparative insight is indispensable not only for the ongoing research within the quantum computing sphere but also for steering the direction of future QECC development.
Navigating through the myriad studies to deduce the optimal QECC for a particular scenario poses a formidable challenge, compounded further when multiple scenarios are considered. Let us consider two hypothetical scenarios: (a) we have a setup with superconducting hardware with qubits where the circuit requiring correction involves just a single qubit and (b) we have a setup with trapped-ion hardware with qubits, and circuit to be corrected includes qubits as well as multi-qubit gates. The optimal QECC for these two situations would differ significantly. The complexity of determining the most appropriate QECC — factoring in the types of qubits, error rates, and other critical parameters — demands a systematic selection strategy. The necessity for a framework that can swiftly identify a suitable QECC based on specific scenario requirements is not just a matter of convenience but a crucial step forward in the practical implementation of QECCs.
@@ I-B Contribution
To our knowledge, this research marks the inaugural effort to establish a unified benchmarking methodology for QECCs. Our objective is to discern how different QECCs are ideally suited for various scenarios and devise a systematic strategy for selecting QECCs appropriately. The contributions of this study can be outlined as follows: Comprehensive Analysis of QECCs: We conduct an in-depth analysis of nine leading QECCs utilizing eight distinct parameters: qubit overhead, threshold, type of error protection, decoding complexity, types of transversal gates, scalability, the range of qubit types it has been realized on, and overall complexity. This enables a detailed comparative overview across different codes. Scenario-dependent QECC Suitability: Our findings emphasize that the selection of the most suitable QECC is contingent upon the specific scenario of its intended application, rather than a theoretical hierarchy of QECC efficacy. We demonstrate this through the use of radar charts, showcasing the trade-offs involved in choosing different QECCs, including variations within a single QECC type. Universal QECC Selection Strategy: We introduce a universal, systematic approach for selecting a QECC tailored to a given scenario. This strategy is designed to guide users in identifying the QECC that best matches their specific requirements, considering the multifaceted nature of quantum computing scenarios. An Adaptive QECC Recommendation Tool: We develop a dynamic tool that assesses user-provided scenario parameters and recommends a list of QECCs, ranked from most to least suitable, along with the maximum distance achievable for each code. This tool can allow for the inclusion of additional QECCs and the swift modification of their parameters without necessitating significant overhauls to the system. This ensures the framework’s adaptability and relevance in the rapidly evolving quantum computing landscape. For instance, should a QECC that has not been previously realized on trapped-ion qubits later achieve this milestone, the framework can easily accommodate this new information.
@@ I-C Paper Structure
Section II provides a brief introduction to QECCs and an extensive examination of the nine QECCs featured in this research. Section III details the eight benchmarking parameters, selection criteria for codes and parameters, and the comparative analysis method. Section IV offers a thorough analysis of the nine QECCs against the selected parameters. Section V discusses a simplified code selection strategy and introduces our QECC recommendation tool. The paper concludes in Section VI.
@@ II-B Descriptions of QECCs Under Study
Shor Code: The Shor Code [26] marks a significant advancement in quantum error correction by being the first to correct arbitrary quantum errors, including both bit-flip and phase-flip errors, in a qubit. It encodes one qubit of information into nine physical qubits through a two-layer structure: the first layer addresses bit-flip errors by tripling the qubit, and the second layer corrects phase-flip errors by further encoding each of these into three qubits. This innovative approach allows the Shor Code to correct any single-qubit error across its nine qubits. It is described as a degenerate code, where different error types can affect the codewords similarly. The two basis states of the code are represented as and , intricately constructed from the superposition and entanglement of the nine qubits. The two basis states for the code are defined as: and . This structure ensures that the Shor Code corrects one X error per trio of qubits and a single Z error across nine qubits, making it a comprehensive quantum error-correcting code. X errors are managed by a correction circuit for each trio, identifying and fixing bit-flips. Z errors are addressed by comparing sign differences between trios with CNOT gates. Its degeneracy allows it to correct a phase flip in any qubit of a trio effectively. Although capable of correcting three bit flips in separate trios, it’s primarily a single-error corrector due to its limitations with multiple errors. Fig. 2 shows the Shor Code’s encoding, error introduction, and correction process.
@@
Steane Code: The Steane Code [35], also known as the 7-qubit code, was introduced as a part of the Calderbank-Shor-Steane (CSS) [27, 28] family of quantum error-correcting codes. It represents a significant advancement in the field of quantum error correction, offering a symmetrical approach to correcting both bit-flip and phase-flip errors. The Steane Code maps one logical qubit to seven physical ones, drawing from the [7,4,3] Hamming code to correct single-bit errors and extending it to fix both bit-flip and phase-flip errors in qubits. Its CSS construction enables independent correction of these errors, streamlining the correction process and simplifying the required quantum circuits. The logical zero and one states in the Steane Code are defined using superpositions of the Hamming code’s even and odd weight codewords, respectively. The Steane Code employs syndrome measurements for error detection, which uniquely indicate the presence and type of error without collapsing the quantum state. This code is comprised of 6 stabilizer generators: , , , , and . Due to this structure, the Steane Code can independently correct X and Z errors using separate circuits. This dual capability is central to its functionality, allowing for the efficient correction of single-qubit errors in quantum systems. The logical and states of this code are: and ; where . Fig. 3 depicts the error-detecting circuit of the 7-qubit Steane code.
@@
Toric Code: The Toric Code [18] represents a novel approach to quantum error correction by leveraging topological properties. Its use of a two-dimensional lattice structure is unique, making it resistant to a broad class of local errors. In the Toric Code, qubits are placed on the edges of a two-dimensional lattice on a torus. The topology of the lattice defines logical qubits and error correction is achieved by measuring vertex and plaquette operators. These operators are defined as: . In these expressions, and represent Pauli X and Z operators applied to the qubits. The products are taken over qubits around a vertex for vertex operators, and around a plaquette for plaquette operators. Errors in the Toric Code manifest as excitations in this lattice structure. Bit-flip errors disrupt vertex operators, while phase-flip errors disrupt plaquette operators. The code detects errors by checking for these excitations. The topological nature of the code allows for the identification and correction of errors based on the collective state of multiple qubits, rather than individual qubit states. Fig. 4 represents the toric code and the acting of the X and Z stabilizers on its unique lattice structure.
@@
Surface Code: Surface Codes [36] are a class of quantum error-correcting codes that extend the principles of the Toric Code to a planar geometry, meaning it is not modeled around a torus. They are known for their high threshold error rate, making them a promising candidate for scalable quantum computation. Surface codes, similar to toric codes in stabilizers and error detection, have two main layouts: unrotated [30] and rotated [37]. Unrotated surface codes use a square lattice, with data qubits on square edges, Z-stabilizers on surfaces, and X-stabilizers at vertices. Rotated surface codes feature a tilted lattice, with X and Z stabilizers alternating in a chessboard pattern. The difference in lattice arrangement impacts their error-correction efficiency, with rotated codes generally preferred for long-distance QEC due to a higher error threshold and simpler implementation. Error detection in Surface Codes is accomplished through the measurement of these stabilizer operators. A non-trivial measurement outcome (eigenvalue -1) indicates an error. The planar nature of Surface Codes facilitates the implementation of fault-tolerant gate operations and scalable quantum computing architectures. Fig. 5 shows unrotated and rotated layouts of a surface code.
@@
Heavy Hexagon (HH) Code: Heavy hexagon codes [41] leverage a hexagonal lattice enhanced with strategically placed ’heavy’ qubits, improving error correction efficiency through increased qubit connectivity. This design boosts error tolerance, crucial for precise syndrome measurements in identifying and correcting bit-flip (X) and phase-flip (Z) errors. The mechanism ensures the integrity of quantum information by accurately pinpointing and fixing errors. This lattice stabilizer subsystem code synthesizes elements from both Bacon-Shor and surface-code frameworks, and encodes a single logical qubit into physical qubits where denotes the code distance. This lattice is characterized by its sparse connectivity, typically not exceeding three connections per qubit, making it a robust candidate for systems with fixed-frequency transmon qubits that are vulnerable to frequency collision errors. On this lattice, data qubits and ancillas are strategically distributed along the hexagonal vertices and edges. A designated subset of these ancilla qubits serves as flag qubits, crucial for the detection of high-weight errors that may result from a limited number of fault events. The stabilizers tasked with the detection of X-type errors are constituted by the product of weight-two Z-type gauge operators, which collectively form the surface code stabilizers. Meanwhile, the X-type stabilizers, reminiscent of columnar operators in the Bacon-Shor code, are deduced by aggregating products of weight-four and weight-two X-type gauge operators. Fig. 8 represents the layout and stabilizer circuits of a heavy-hexagon code.
Bivariate Bicycle (BB) Code: This code [42], is a significant advancement in the field of quantum error correction by IBM. This innovative code belongs to the family of Abelian 2BGA codes known for their efficiency in syndrome measurement within quantum computing, particularly due to their time-optimal circuits that fit neatly into a two-layered setup. This arrangement extends the square-lattice design, optimizing it further for use with surface codes.
Unlike traditional 2D grid structures, the BB code’s qubit connectivity graph unfolds into two planar subgraphs, each with a degree of three, offering a more complex interconnectivity. It utilizes and check operators, each with a weight of six. The syndrome measurement circuit is shown in Fig. 9. Including the ancilla qubits necessary for operation, the BB code achieves an encoding efficiency that exceeds that of the conventional surface code. Specifically, a BB Code described by parameters necessitates ancilla qubits for encoding, resulting in an effective ancilla-adjusted encoding rate of . A standout among BB Codes is the Gross Code, , a specific type of BB QLDPC code, which is remarkable for requiring fewer physical and ancilla qubits for syndrome extraction compared to the surface code, maintaining the same logical qubit count and distance. The term ”Gross” here cleverly alludes to the code’s efficiency, as ”gross” traditionally means a dozen dozens, highlighting the code’s compactness with an ancilla-adjusted rate of . For comparison, the surface code with a distance of exhibits a much lower ancilla-adjusted rate of , underscoring the BB code’s superior efficiency.
@@ III-A Explanation of Benchmarking Parameters
Qubit Overhead: Qubit overhead is a fundamental parameter in assessing the efficiency of a QECC. It refers to the number of physical qubits needed to encode a logical qubit; lower overhead is preferable as it implies fewer resources for error correction. However, this must be balanced against the code’s ability to correct errors, which often requires more qubits for increased distance. Therefore, analyzing the trade-off between qubit overhead and error correction capabilities is crucial in evaluating the practicality of a QECC. Error Threshold: The error threshold is a critical benchmark for the robustness of a QECC as it represents the maximum physical error rate under which a QECC can effectively correct errors. A higher threshold indicates a more fault-tolerant code, making it more desirable for practical quantum computing where error rates are non-negligible. Error Protection: Error protection capability is the measure of a QECC’s ability to identify and correct a range of quantum errors, including bit flips, phase flips, and their combinations (Pauli errors). The ultimate goal is to design QECCs capable of detecting and correcting all possible quantum errors. A QECC’s versatility in handling diverse error types is a significant indicator of its effectiveness and practical utility in quantum computing environments. Decoding: The availability and efficiency of decoding algorithms are essential for the practical implementation of QECCs. Decoding is the process of interpreting syndrome measurements to identify and correct errors. The more decoding algorithms a QECC is compatible with, the more flexible and adaptable it is. The performance of a QECC is heavily reliant on the efficiency of these algorithms, which should be fast and accurate to maintain the integrity of quantum information. Transversal Gates: Transversal gates play a crucial role in the implementation of quantum error correction. They facilitate operations on encoded qubits without decoding them, thereby preserving the error correction capability. The complexity and type of transversal gates that a QECC supports can significantly affect its functionality and integration with quantum circuits. Advanced techniques like lattice surgery, while complex, offer enhanced capabilities in error correction and logical operations, marking an important consideration in QECC design. Scalability: Scalability is an essential factor for the practical application of QECCs in large-scale quantum computing. It refers to a QECC’s ability to maintain its error correction effectiveness as the system size increases. Scalability is crucial for the advancement of quantum computing, as it determines whether a QECC can be effectively used in more complex and realistic quantum systems. Realization: The realization of a QECC across different qubit types is a testament to its adaptability and practical relevance. The versatility of a QECC in being implemented with various qubit technologies (such as superconducting qubits, trapped-ions, etc.) is vital for its applicability in diverse quantum computing architectures. This parameter is a strong indicator of a QECC’s potential for widespread adoption in the quantum computing field. Complexity: The structural and operational complexity of a QECC impacts its implementation ability and efficiency. This encompasses the intricacy of its encoding, error detection, and correction processes. A less complex QECC might be easier to implement but may offer limited error correction capabilities. Conversely, a more complex QECC might provide robust error correction but at the cost of increased resource demands and operational challenges. Balancing complexity with effectiveness and resource efficiency is key in the design and selection of QECCs for practical quantum computing.